π-π and CH-π DISPERSION INTERACTIONS IN THE N_{2}-C_{5}H_{5}N AND CH_{4}-C_{5}H_{5}N DIMERS

\underline{CHAKREE TANJAROON}^{a} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, T6G 2G2, Canada.

We will discuss the results of our high resolution spectroscopic studies of the dinitrogen-pyridine (N_{2}-C_{5}H_{5}N) and methanepyridine (CH_{4}-C_{5}H_{5}N) dimers. The two dimers represent simple binary van der Waals systems that are good prototypes for studying the weak π-π and CH-π bonds involving polar and non-polar aromatic molcules. The pyridine molecule is of great interest because many of its derivatives are prevalently found in plants and microorganisms. Our preliminary results show that the dimers adopt T-shaped configurations. The N_{2} lies perpendicularly to the aromatic plane with its center of mass closes to the c-inertial axis of the free pyridine molecule; similarly, the CH_{4} molecule in CH_{4}-C_{5}H_{5}N sits above the aromatic plane. These are the only configurations presently observed for both dimers in our experiments. The talk will focus on the rotational spectra and structural properties of these dimers.

aSupported by an Alberta Ingenuity Postdoctoral Fellowship.